Simplify the exponential expression.
(.25xy)(-20x^-1y^-5)

Answer by checkley71(8403)

(.25xy)(-20x^-1y^-5)
.25*-20=-5
x^1*-x^-1=x^0=1
y^1*y^-5=y^-4=1/y^4
so we have:
-5/y^4 answer.